Louisiana State University Health Sciences Center Administration & Finance
 

Friday, May 05, 2006

  • CSRB switch gear: The search for the cause(s) of the power outages discovered three damaged electrical feeders that serve the building. Plans are underway to replace the most critical feeder during the evening of Tuesday, May 9th and the other two feeders on Saturday, May 13th. Specifics on the scheduled building shut downs will be published via mass email shortly.
  • Temporary repairs at the Dental School continue.
    • The contractors generator has been delivered to the site.
    • The transfer switch and two tap boxes needed to connect power to two of the buildings risers have also been received. The contractor will be installing the tap boxes over the weekend.
    • The installation of new conduit will also be finished this weekend.
    • Environmental remediation in the basement continues; it is approximately 90% complete.
    • All of the pipe insulation in the basement is complete. The temporary chillers are to be operating again on Saturday.
  • The contractor awarded the 3rd environmental remediation contract on the downtown campus mobilized today and work is underway on that contract.
  • The fire alarm contractor continued their work in the Resource Center. They are running conduit and pulling wire to re-route wires that were originally run through the slab.